OpenStudy (anonymous):

PLEASE HELP ME I WILL GIVE MEDAL AND BE SO THANKFUL! 4. Solve for x. 2x^2-7x=4 A) {-3/2,4} B) {-3,4} C) {-1/2,4} D) {-3/2,-4} 5.Solve for x. (-12x^3y^4+6x^2y^2)divided by 6x^2y^2 A) -2xy^2+3y+1 B) 3y-2xy C) 2x^2y-1 D) -2xy^2+1 6. Solve for c. A=1/2(a+b+c). A) c=2A-(a-b) B) c=2A+a+b C) c=2A-a-b D) c=2Aab

OpenStudy (anonymous):

@Here_to_Help15 help? Lol

OpenStudy (anonymous):

Anyone? Pls?

OpenStudy (here_to_help15):

hmmm im still mad at u for not calling me first :P

OpenStudy (anonymous):

Hahaha I'm sorryyy please forgive me :)

OpenStudy (here_to_help15):

i forgive u : P

OpenStudy (anonymous):

Yyyyaaaaaayyyyyyyy! Thank you! :) now can you help? :)

OpenStudy (lochana):

4.C

OpenStudy (anonymous):

Oh thank you @lochana :)

OpenStudy (here_to_help15):

friggin lochana lol

OpenStudy (anonymous):

Hahahahahaha

OpenStudy (midhun.madhu1987):

4. C) {-1/2,4}

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

Do you guys think you could help with 5 and 6 too? Please :)

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

4. http://www.wolframalpha.com/input/?i=2x%5E2-7x-4%3D0

OpenStudy (nurali):

5. -2xy^2+3y+1

OpenStudy (anonymous):

Oh my god thank you lol @Nurali

OpenStudy (nurali):

6. A=1/2(a+b+c). 2A=a+b+c 2A-a-b=c c=2A-a-b

OpenStudy (nurali):

Anytime.

OpenStudy (anonymous):

Wait @Nurali is number 6 answer c=2A-a-b?

OpenStudy (nurali):

Yes

OpenStudy (anonymous):

Oh okay thank you!!!! <3

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

I already gave it to someone but I will open a new question so I can give you one